If y = 0.02 sin (30x − 400t) (SI units),

the frequency of the wave is

30 Hz
b. 15/ Hz
c. 200/ Hz
d. 400 Hz
e. 800 Hz

the angular frequency of the wave is

a. 30 rad/s
b. 30/2 rad/s
c. 400/2 rad/s
d. 400 rad/s
e. 40/3 rad/s

the velocity of the wave is

a. 3/40 m/s
b. 40/3 m/s
c. 60/400 m/s
d. 400/60 m/s
e. 400 m/s
Answer #1

Y = A sin (kx - wt)

a) f = w / 2 pi = 400 / 2 pi

= 200 / pi [option c]

b) w = 2 pi f

w = 400 rad/s [option d]

c) v = w / k = 400 / 30 = 40 / 3 m/s

40/3 m/s [option b]